OpenStudy (anonymous):

Part 1 [2 points] Is the following equation true? Part 2 [6 points] Use complete sentences to explain the properties used in making your decision. 3 log 2 (x + 1/2) log 2 (y - 3) log 2 (z) = log 2 (x^3√y/z^3)

OpenStudy (anonymous):

I'm so confused :(

OpenStudy (perl):

im afraid you didnt copy it correctly

OpenStudy (anonymous):

ill draw the equation

OpenStudy (perl):

thanks

OpenStudy (anonymous):

\[3\log_{2} x+\frac{ 1 }{ 2 } \log_{2} y-3 \log_{2} z = \log_{2} \frac{ x^3\sqrt{y} }{ z^3 }\]

OpenStudy (perl):

yes its true, by laws of logs

OpenStudy (anonymous):

How would I explain the properties?

OpenStudy (perl):